# Locally Repairable Convertible Codes: Improved Lower Bound and General Construction

### v2

@@ IV A General Construction of LRCCs
    For an optimal $g(r+\delta-1),\ell r,d]_{q}-$LRC $\mathcal{C}$ with $(r,\delta)_{a}$-locality and achieving the minimum Hamming distance bound in (3), the $\mathcal{C}$ admits a parity check matrix $H$ of the form (1) where $A_{1},A_{2},\dots,A_{g}$ are $(\delta-1)\times(r+\delta-1)$ matrices over $\mathbb{F}_{q}$, and $B_{1},B_{2},\dots,B_{g}$ are $((g-\ell)r)\times(r+\delta-1)$ matrices over $\mathbb{F}_{q}$.
@@
$H||_{\mathcal{P}_{1}\cup\mathcal{P}}=\left(\begin{array}[]{ccc|ccc|ccc}A_{1}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}\\ \vdots&\ddots&\vdots&\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}\\ \bm{0}&\cdots&A_{s}&\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}\\ \hline\cr\bm{0}&\cdots&\bm{0}&A_{\ell+1}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&A_{\ell+h}&\bm{0}&\cdots&\bm{0}\\ \hline\cr\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}&A_{\ell+h+1}&\cdots&\bm{0}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&A_{g}\\ B_{1,1}&\cdots&B_{s,1}&B_{\ell+1,1}&\cdots&B_{\ell+h,1}&B_{\ell+h+1,1}&\cdots&B_{g,1}\\ \hline\cr B_{1,2}&\cdots&{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B_{s,2}}&B_{\ell+1,2}&\cdots&B_{\ell+h,2}&B_{\ell+h+1,2}&\cdots&B_{g,2}\end{array}\right)\triangleq\left(\begin{array}[]{c|c|c}W_{11}&\bm{0}&\bm{0}\\ \hline\cr\bm{0}&W_{12}&\bm{0}\\ \hline\cr M_{11}&M_{12}&M_{13}\\ \hline\cr N_{11}&N_{12}&N_{13}\end{array}\right),$ (25)
@@
$\begin{pmatrix}I&\bm{0}&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}&\bm{0}\\ \bm{0}&\bm{0}&I&\bm{0}\\ \bm{0}&\bm{0}&-N_{13}M_{13}^{-1}&I\end{pmatrix}\cdot\begin{pmatrix}W_{11}&\bm{0}&\bm{0}\\ \bm{0}&W_{12}&\bm{0}\\ M_{11}&M_{12}&M_{13}\\ N_{11}&N_{12}&N_{13}\end{pmatrix}=\begin{pmatrix}W_{11}&\bm{0}&\bm{0}\\ \bm{0}&W_{12}&\bm{0}\\ M_{11}&M_{12}&M_{13}\\ N_{11}-N_{13}M_{13}^{-1}M_{11}&N_{12}-N_{13}M_{13}^{-1}M_{12}&\bm{0}\end{pmatrix}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\triangleq H_{1}^{*}}.$ (26)
@@
Since every $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[t]$ has the same part $H||_{\mathcal{P}}$, thus the punctured code $\mathcal{C}_{i}|_{[(s+h)(r+\delta-1)]}$ from $\mathcal{C}_{i}$ comes from the same elementary row operations, i.e., each $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ is left-multiplied by the first matrix in the left-hand side of (26). Then, $\mathcal{C}_{i}|_{[(s+h)(r+\delta-1)]}$ has the parity check matrix $\overline{H}_{i}$ for $i\in[t]$, i.e.,
@@
and the code symbols $\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]}$ correspond to the read symbols for $\bm{c}_{i}$ for $1\leqslant i\leqslant t$. The final code $\mathcal{C}_{F}$ is defined as
@@
by (3). Therefore, by Lemma 1, we have $\dim(\tilde{\mathcal{C}})=\ell r$ since $(g-\ell-h)(r+\delta-1)<d$, which together with (32) show that $\mathcal{C}_{F}=\tilde{\mathcal{C}}$.
@@
Verify that the access cost is optimal: Combining (28) and (29), we derive that the write access cost is $h(r+\delta-1)$ and the read access cost is thrthr. According to Corollary 1, we have that
$=t(\delta+h(r+\delta-1)+(t-1)sr-1+(s(t-1)-1)(\delta-1))$
$\quad-(t-1)(\ell+h)(r+\delta-1)$
$=h(r+\delta-1),$
and
$=(\ell+h)(r+\delta-1)-2(\delta+h(r+\delta-1))$
$\quad-(t-1)sr+2-(s(t-1)-1)(\delta-1)$
$=(s-h-1)(r+\delta-1)+r$
$>0,$
where the last inequality holds by $s=\ell/t>g-\ell\geqslant h$. Therefore,
$\rho_{r}$ $\geqslant t\left(sr-\Delta+(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor\right)$
$=t(sr+h(r+\delta-1)-s(r+\delta-1)+\delta-1+(\delta-1)(s-h-1))$
$=thr.$
The read and write access costs both achieve the lower bound in Corollary 1. Thus, the LRCC above is access-optimal, which completes the proof. ∎
In fact, the proof of Theorem 3 gives the following construction of access-optimal LRCCs.
@@
Since $\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{t}$ are the same code with parity check matrix $H||_{\mathcal{P}_{1}\cup\mathcal{P}}$, then modify the conversion procedure function (28) to
@@
which shows (30) and (31) also hold. Moreover, the equations (30) and (31) hold in this case. Therefore, the new LRCC is access-optimal, and the initial and final codes are optimal LRCs, where the initial codes are equal.
@@ V-A Super-linear length
Base Code A ([28, Construction A]): Let $d\geqslant\delta+1$ and $n=m(r+\delta-1)$, where $d,\delta,r,m$ are positive integers. For $i\in[m]$, let $G_{i}=\{g_{i,1},g_{i,2},\dots,g_{i,r+\delta-1}\}$ be a $(r+\delta-1)$-subset of $\mathbb{F}_{q}$. Then, for each $i\in[m]$, we can construct a $(d-1)\times(r+\delta-1)$ Vandermonde matrix with generating set $G_{i}$ of the form $(U_{i}^{\top},V_{i}^{\top})^{\top}$, where
@@
Moreover, the initial and final codes have super-linear length and both are optimal LRCs.
@@
In fact, $\mathcal{P}_{i}=\{(i-1)u+1,\dots,iu\}$ for $i\in[\frac{m-1}{u}]$, and $\mathcal{P}=\{m\}$. According to Corollary 2, the LRCC given by Construction A is access-optimal. According to Lemma 11, the initial codes are optimal $(u+1)(r+\delta-1),ur,r+\delta]_{q}-$LRC with $(r,\delta)_{a}$-locality.
@@
We give an example for access-optimal $(8,1)_{7^{2}}$ LRCC with parameters $u=2,r=2,\delta=2,d=4,m=17$ in Theorem 4. Since $x^{2}-3$ is an irreducible polynomial over $\mathbb{F}_{7}$, let $\mathbb{F}_{7^{2}}=\mathbb{F}_{7}[x]/(x^{2}-3)=\mathbb{F}_{7}(\beta)$, where $\beta^{2}-3=0$. Then, $\{1,\beta\}$ is a basis of $\mathbb{F}_{7^{2}}$ over $\mathbb{F}_{7}$. Clearly, $\mathbb{F}_{7}^{*}$ is a multiplicative subgroup of $\mathbb{F}_{7^{2}}^{*}$. There are $8$ cosets of $\mathbb{F}_{7}^{*}$, i.e., $\mathbb{F}_{7}^{*}$ and $(i+\beta)\mathbb{F}_{7}^{*},i=0,1,\dots,6$. Let $G_{0}=\{5,6,0\},G_{1}=\{1,2,3\},G_{2}=\{3,4,5\}$. Obviously, $G_{0},G_{1},G_{2}$ satisfy the relationship (37). Therefore, the corresponding initial code $\mathcal{C}_{0}$, based on Base Code A and $G_{0},G_{1},G_{2}$, has the code length $9$ which is large than the size of $\mathbb{F}_{7}$. According to Remark 7, let the initial code $\mathcal{C}_{i}$ is given by Base Code A and $(i-1+\beta)G_{0},(i-1+\beta)G_{1},(i-1+\beta)G_{2}$ for $i\in\{1,2,3,\dots,7\}$. Then, the final code $\mathcal{C}_{F}$, based on Base Code A and $G_{0},G_{1},G_{2},(i-1+\beta)G_{1}\cup(i-1+\beta)G_{2},i=1,2,\dots,7$, has the code length $17\times 3=51>49$. It is easy to check that $\mathcal{C}_{F}$ satisfies the requirement (37). Therefore, the initial and final codes are both with super-linear length. Also, the initial codes $\mathcal{C}_{0},\mathcal{C}_{1},\dots,\mathcal{C}_{7}$ are the same one $\mathcal{C}_{0}$.
@@ V-B Maximally recoverable codes
Base Code B([4, Construction A]): Let $\alpha_{1},\alpha_{2},\dots,\alpha_{r+\delta-1}$ be $r+\delta-1$ distinct elements in $\mathbb{F}_{q}^{*}$. Let
@@
Indeed, $\mathcal{P}_{i}=\{(i-1)u+1,\dots,iu\}$ for $i\in[\frac{m-h/r}{u}]$, and $\mathcal{P}=\{m-h/r+1,\dots,m\}$. This implies that the initial code $\mathcal{C}_{i}$ has parity check matrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[t]$ and the final code $\mathcal{C}_{F}$ is by means of the parity check matrix $H$, where $H$ is given by (38). Based on Lemma 12 and $r\mid h$, the minimum Hamming distance of $\mathcal{C}$ is $d=\delta+\frac{h}{r}(r+\delta-1)$ which achieves the lower bound (3). According to Theorem 3 where $g=m$ and $\ell=m-h/r$, the LRCC given by Construction A is access-optimal. Noting that $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[t]$ also maintain the structure of Base Code B, therefore, by Lemma 12, the initial codes are $((u+h/r)(r+\delta-1),r,h,\delta,q^{h})$-MR codes and the final code is a $(m(r+\delta-1),r,h,\delta,q^{h})$-MR code. ∎
@@
The $t$ initial codes in Theorem 5 could be the distinct codes. In the following case, the initial codes may be the same one. When $H||_{\mathcal{P}_{i}}=D_{i}\cdot H||_{\mathcal{P}_{1}}$ where
@@
the $t$ initial codes can be the same code by Remark 4 when $h=r$.
@@
The final code given by the conversion procedure function (35) has the parity check matrix
$H=\left(\begin{array}[]{cccccc|c}1&1&1&0&0&0&\bm{0}_{1\times 9}\\ 0&0&0&1&1&1&\bm{0}_{1\times 9}\\ \hline\cr\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\\ 1+\beta&1+2^{2}\beta&1+3^{2}\beta&1+\beta&1+2^{2}\beta&1+3^{2}\beta&H_{2}\\ (1+\beta)^{7}&(1+2^{2}\beta)^{7}&(1+3^{2}\beta)^{7}&(1+\beta)^{7}\gamma&(1+2^{2}\beta)^{7}\gamma&(1+3^{2}\beta)^{7}\gamma&\end{array}\right),$
where $H_{1}||_{\{1,2\}}=H||_{\{1,2\}},H_{2}||_{\{1,2\}}=H||_{\{3,4\}}$. Clearly, $\mathcal{C}_{F}$ is an MR code.

### v3

@@ IV A General Construction of LRCCs
    For an optimal $g(r+\delta-1),\ell r,d]_{q}-$LRC $\mathcal{C}$ with $(r,\delta)_{a}$-locality and achieving the minimum Hamming distance bound in (3), the $\mathcal{C}$ admits a parity check matrix $H$ of the form (1), where $A_{1},A_{2},\dots,A_{g}$ are $(\delta-1)\times(r+\delta-1)$ matrices over $\mathbb{F}_{q}$, and $B_{1},B_{2},\dots,B_{g}$ are $((g-\ell)r)\times(r+\delta-1)$ matrices over $\mathbb{F}_{q}$.
@@
$H||_{\mathcal{P}_{1}\cup\mathcal{P}}=\left(\begin{array}[]{ccc|ccc|ccc}A_{1}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}\\ \vdots&\ddots&\vdots&\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}\\ \bm{0}&\cdots&A_{s}&\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}\\ \hline\cr\bm{0}&\cdots&\bm{0}&A_{\ell+1}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&A_{\ell+h}&\bm{0}&\cdots&\bm{0}\\ \hline\cr\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}&A_{\ell+h+1}&\cdots&\bm{0}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&A_{g}\\ B_{1,1}&\cdots&B_{s,1}&B_{\ell+1,1}&\cdots&B_{\ell+h,1}&B_{\ell+h+1,1}&\cdots&B_{g,1}\\ \hline\cr B_{1,2}&\cdots&B_{s,2}&B_{\ell+1,2}&\cdots&B_{\ell+h,2}&B_{\ell+h+1,2}&\cdots&B_{g,2}\end{array}\right)\triangleq\left(\begin{array}[]{c|c|c}W_{11}&\bm{0}&\bm{0}\\ \hline\cr\bm{0}&W_{12}&\bm{0}\\ \hline\cr M_{11}&M_{12}&M_{13}\\ \hline\cr N_{11}&N_{12}&N_{13}\end{array}\right),$ (25)
@@
$\begin{pmatrix}I&\bm{0}&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}&\bm{0}\\ \bm{0}&\bm{0}&I&\bm{0}\\ \bm{0}&\bm{0}&-N_{13}M_{13}^{-1}&I\end{pmatrix}\cdot\begin{pmatrix}W_{11}&\bm{0}&\bm{0}\\ \bm{0}&W_{12}&\bm{0}\\ M_{11}&M_{12}&M_{13}\\ N_{11}&N_{12}&N_{13}\end{pmatrix}=\begin{pmatrix}W_{11}&\bm{0}&\bm{0}\\ \bm{0}&W_{12}&\bm{0}\\ M_{11}&M_{12}&M_{13}\\ N_{11}-N_{13}M_{13}^{-1}M_{11}&N_{12}-N_{13}M_{13}^{-1}M_{12}&\bm{0}\end{pmatrix}\triangleq H_{1}^{*}.$ (26)
@@
Since all the matrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[t]$ share the common submatrix $H||_{\mathcal{P}}$, thus the punctured code $\mathcal{C}_{i}|_{[(s+h)(r+\delta-1)]}$ from $\mathcal{C}_{i}$ comes from the same elementary row operations, i.e., each $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ is left-multiplied by the first matrix in the left-hand side of (26). Then, $\mathcal{C}_{i}|_{[(s+h)(r+\delta-1)]}$ has the parity check matrix $\overline{H}_{i}$ for $i\in[t]$, i.e.,
@@
and the code symbols $\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]}$ correspond to the read symbols from $\bm{c}_{i}$ for $1\leqslant i\leqslant t$. The final code $\mathcal{C}_{F}$ is defined as
@@
by (3). Therefore, by Lemma 1, we have $\dim(\tilde{\mathcal{C}})=\ell r$ since $(g-\ell-h)(r+\delta-1)<d$, which together with (32) shows that $\mathcal{C}_{F}=\tilde{\mathcal{C}}$.
@@
Verify that the access cost is optimal: Combining (28) and (29), we derive that the write access cost is $h(r+\delta-1)$ and the read access cost is thrthr. According to Corollary 2, we have that $\rho_{w}\geqslant h(r+\delta-1)$ and $\rho_{r}\geqslant thr$. Thus, the read and write access costs both achieve the lower bounds in Corollary 2. Thus, the LRCC above is access-optimal, which completes the proof. ∎
In fact, the proof of Theorem 3 gives the following general construction of access-optimal LRCCs.
@@
Since $\mathcal{C}_{1}=\mathcal{C}_{2}=\dots=\mathcal{C}_{t}$ with parity check matrix $H||_{\mathcal{P}_{1}\cup\mathcal{P}}$, then modify the conversion procedure function (28) to
@@
which shows (30) and (31) also hold. Therefore, the new LRCC is access-optimal, and the initial and final codes are optimal LRCs, where the initial codes are equal.
@@ V-A Super-linear length
Base Code A ([48, Construction A]): Let $d\geqslant\delta+1$ and $n=m(r+\delta-1)$, where $d,\delta,r,m$ are positive integers. For $i\in[m]$, let $G_{i}=\{g_{i,1},g_{i,2},\dots,g_{i,r+\delta-1}\}$ be a $(r+\delta-1)$-subset of $\mathbb{F}_{q}$. Then, for each $i\in[m]$, we can construct a $(d-1)\times(r+\delta-1)$ Vandermonde matrix with generating set $G_{i}$ of the form $(U_{i}^{\top},V_{i}^{\top})^{\top}$, where
@@
Moreover, the initial and final codes have super-linear length and both are optimal LRCs with respect to the bound in Lemma 3.
@@
In fact, $\mathcal{P}_{i}=\{(i-1)u+1,\dots,iu\}$ for $i\in[\frac{m-1}{u}]$, and $\mathcal{P}=\{m\}$. Based on (28) and (29), the read access cost is $\rho_{r}=\frac{(m-1)r}{u}$ and the write access cost is $\rho_{w}=r+\delta-1$. The LRCC given by Construction A is access-optimal, which corresponds to the case $h=1$ and $t=\frac{m-1}{u}$ for Corollary 3. According to Lemma 11, the initial codes are optimal $(u+1)(r+\delta-1),ur,r+\delta]_{q}-$LRC with $(r,\delta)_{a}$-locality.
@@
We give an example for access-optimal $(8,1)_{7^{2}}$ LRCC with parameters $u=2,r=2,\delta=2,d=4,m=17$ in Theorem 4. Since $x^{2}-3$ is an irreducible polynomial over $\mathbb{F}_{7}$, let $\mathbb{F}_{7^{2}}=\mathbb{F}_{7}[x]/(x^{2}-3)=\mathbb{F}_{7}(\beta)$, where $\beta^{2}-3=0$. Then, $\{1,\beta\}$ is a basis of $\mathbb{F}_{7^{2}}$ over $\mathbb{F}_{7}$. Clearly, $\mathbb{F}_{7}^{*}$ is a multiplicative subgroup of $\mathbb{F}_{7^{2}}^{*}$. There are $8$ cosets of $\mathbb{F}_{7}^{*}$, i.e., $\mathbb{F}_{7}^{*}$ and $(i+\beta)\mathbb{F}_{7}^{*},i=0,1,\dots,6$. Let $G_{0}=\{5,6,0\},G_{1}=\{1,2,3\},G_{2}=\{3,4,5\}$. Obviously, $G_{0},G_{1},G_{2}$ satisfy the relationship (37). Therefore, the corresponding initial code $\mathcal{C}_{0}$, based on Base Code A and $G_{0},G_{1},G_{2}$, has the code length $9$ which is larger than the size of $\mathbb{F}_{7}$. According to Remark 8, let the initial code $\mathcal{C}_{i}$ be given by Base Code A and $(i-1+\beta)G_{0},(i-1+\beta)G_{1},(i-1+\beta)G_{2}$ for $i\in\{1,2,3,\dots,7\}$. Then, the final code $\mathcal{C}_{F}$, based on Base Code A and $G_{0},G_{1},G_{2},(i-1+\beta)G_{1}\cup(i-1+\beta)G_{2},i=1,2,\dots,7$, has the code length $(2\times 8+1)\times 3=51>49$, since there are $2$ groups of samples from each of $8$ initial codes and there is one group of samples corresponding to the global parity check symbols. It is easy to check that $\mathcal{C}_{F}$ satisfies the requirement (37). Therefore, the lengths of initial and final codes are both beyond the field size, which can not constructed by Tamo-Barg codes [19]. Also, the initial codes are the same one, i.e., $\mathcal{C}_{0}=\mathcal{C}_{1}=\dots=\mathcal{C}_{7}$.
@@ V-B Maximally recoverable codes
Base Code B([55, Construction A]): Let $\alpha_{1},\alpha_{2},\dots,\alpha_{r+\delta-1}$ be $r+\delta-1$ distinct elements in $\mathbb{F}_{q}^{*}$. Let
@@
Recall that $\mathcal{P}_{i}=\{(i-1)u+1,\dots,iu\}$ for $i\in[\frac{m-h/r}{u}]$, and $\mathcal{P}=\{m-h/r+1,\dots,m\}$. This implies that the initial code $\mathcal{C}_{i}$ has parity check matrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[t]$ and the final code $\mathcal{C}_{F}$ is by means of the parity check matrix $H$, where $H$ is given by (38). Based on Lemma 12 and $r\mid h$, the minimum Hamming distance of $\mathcal{C}$ is $d=\delta+\frac{h}{r}(r+\delta-1)$ which achieves the lower bound (3). According to Theorem 3 where $g=m$ and $\ell=m-h/r$, the LRCC given by Construction A is access-optimal. Noting that $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[t]$ also maintain the structure of Base Code B, therefore, by Lemma 12, the initial codes are $((u+h/r)(r+\delta-1),r,h,\delta,q^{h})$-MR codes and the final code is an $(m(r+\delta-1),r,h,\delta,q^{h})$-MR code. ∎
@@
In general, the $t$ initial codes in Theorem 5 could be the distinct codes. In the following case, the initial codes may be the same one. When $H||_{\mathcal{P}_{i}}=D_{i}\cdot H||_{\mathcal{P}_{1}}$, where
@@
the $t$ initial codes can be the same code by Remark 5 provided that $h=r$.
@@
The final code $\mathcal{C}_{F}$ given by the conversion procedure function (35) has the parity check matrix
$H=\left(\begin{array}[]{cccccc|c}1&1&1&0&0&0&\bm{0}_{1\times 9}\\ 0&0&0&1&1&1&\bm{0}_{1\times 9}\\ \hline\cr\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\\ 1+\beta&1+2^{2}\beta&1+3^{2}\beta&1+\beta&1+2^{2}\beta&1+3^{2}\beta&H_{2}\\ (1+\beta)^{7}&(1+2^{2}\beta)^{7}&(1+3^{2}\beta)^{7}&(1+\beta)^{7}\gamma&(1+2^{2}\beta)^{7}\gamma&(1+3^{2}\beta)^{7}\gamma&\end{array}\right).$
It is easy to verify that $\mathcal{C}_{F}$ is an MR code.